# Towards autonomous photogrammetric forest inventory using a lightweight under-canopy robotic drone

### v1

@@ 4.2.1 Photogrammetric processing
Results of the photogrammetric processing are given in Table 5. In the image alignment process, all images were aligned in each dataset, which indicated successful processing. Re-projection errors varied between 0.29 and 0.39 pixels. The estimated flight GSDs (Ground Sampling Distance) ranged between 4.08 and 6.17 mm at ground level. Distances from the trees varied between 0.55 and 7.94 m, resulting in GSDs of 0.85 - 12.34 mm at the objects of interest. Point clouds from datasets II-IV had similar point densities of 2.63-2.84 points/cm2, while dataset I provided a slightly higher point density of 3.58 points/cm2.
Results of accuracy estimation using Check Points (CPs) are shown in Table 6. The cases without GCPs or with one GCP at the beginning of the trajectory provided similar 2D and 3D RMSEs, of 11.36-11.72 cm and 11.37-11.73 cm, respectively. When two GCPs were used, the 2D and 3D RMSE improved to 5.26 cm and 5.48 cm, respectively. Adding an additional GCP resulted in a significant improvement in the 2D and 3D RMSE, reducing them to 1.41 cm and 1.48 cm, respectively.
Table 5: Photogrammetric processing results. Img. count: Number images/Number of aligned images; Number of GCPs; Completeness % (Number of trees found/Number of total trees); Reprojection error (pix); Number of tie points (in millions); Point density (points(cm2) in ”Evo-medium” test site. I: Trajectory 1, II: Trajectory 2, III: Trajectory 3, IV: Trajectory 4, Combined: Trajectories 1-4 merged, and for Trajectory 1 I0GCP: no GCPs, I1GCP: 1 GCP, I2GCP: 2 GCPs, I3GCP: 3 GCPs
@@
Table 6: Accuracy estimation (Root Mean Square Error; RMSE) using Check Points (CPs) and Camera Location in ”Evo-medium” test site, for trajectory 1: I0GCP: no GCPs, I1GCP: 1 GCP, I2GCP: 2 GCPs, I3GCP: 3 GCPs
@@ 4.2.2 Tree detection
Figure 12 shows post-processed trajectories for dataset I-IV, as well as reference and detected trees, and the boundaries of the point clouds. I and II followed almost the same trajectory, while III and IV had similar trajectories. Dataset I trajectory was smoothest and it did not include any sharp turns. Other 3 trajectories had one or more sharper turns. For datasets I, II, III and IV, the total number of reference trees inside the point cloud boundary was 29, 29, 31 and 32, of which 23, 18, 19 and 20 were found from drone-collected point cloud. The resulting completeness values were 79.31%, 62.07%, 61.29% and 62.5%, respectively. The total number of different trees found was 29 out of 33 (87,88%) from dataset I-IV. However, when all the flights were processed simultaneously (combined) that resulted only $3$ completeness, where only 12 of 36 trees was found. When utilizing 0, 1, 2 and 3 gcps in dataset I completeness values were 72.41 %, 68.97 %, 75.86 % and 79.31 %, respectively (Table 5).
In dataset I, the stem detection algorithm failed to detect one large tree and two small trees located near the drone’s trajectory, despite these trees being well-reconstructed in the point cloud. Similarly, number of well-reconstructed trees but not detected by algorithm was 4, 5 and 5, in dataset II, III and IV, respectively. The three extra trees that were not identified in dataset I, were situated further from the trajectory and exhibited poor reconstruction quality in the point cloud. The datasets II trough IV each contained 7 poorly reconstructed trees. Many of these trees were located at a greater distance from the trajectory, while others were obscured from the camera’s view by intervening trees.
@@ 4.2.3 DBH estimation
RMSE and bias of the DBH estimates of datasets I-IV were similar, ranging from 3.33 to 3.86 cm (10.69-12.98%) and -0.67 to 1.40 (-2.11-4.71%), respectively (Table 7). Processing all the flights simultaneously (Combined) provided only marginally better results with an RMSE of 2.91 cm (9.14%) and a bias of 1.51 cm (4.74%).
When considering the impact of the number of GCPs, the utilization of 3, 2, 1 and 0 GCPs in photogrammetric processing yielded slightly poorer results compared to those obtained with all GCPs. The RMSE and bias values ranged from 3.93 to 4.37 cm (13.68%-14.66%) and 1.12 to 2.13 cm (3.89-7.55%), respectively (Table 7).
Tree size had an impact on the DBH estimation accuracy (Table 7). For small and medium-sized trees, with a DBH smaller than 30 cm, DBH estimations achieved superior performance in all dataset, with RMSE and bias varying from 1.16 to 2.75 cm (5.74-12.47%) and -0.01 to 0.62 cm (-0.06-3.14%), respectively. Significantly poorer performance was obtained for trees with a DBH greater than 30 cm, with RMSE and bias ranging from 4.04 to 11.08 cm (8.99-25.44%) and -0.66 to 2.08 cm (-1.53-5.65%), respectively. This is demonstrated in Figure 14, illustrating a positive correlation between tree diameter and DBH estimation error. Notably, when the tree diameter exceeded 30 cm, the DBH estimation errors became more pronounced.
Analysis of the impact of the camera distance from trees did not reveal significant correlation between tree distance and DBH estimation errors (Figure 13).
Figure 12: Post-processed trajectories, reference trees and found trees in Dataset (a) I (b) II (c) III (d) IV..
Figure 13: Scatter plots of DBH errors with respect to distance to trees. Dataset (a) I (b) II (c) III (d) IV.
Figure 14: Scatter plots of DBH errors with respect to diameter of trees. Dataset (a) I (b) II (c) III (d) IV
Table 7: Accuracy of the stem diameter breast height (DBH) in ”Evo-medium” test site. Dataset is split in 3 categories: All, DBH $<$ 30 cm and DBH $>$ 30 cm trees. Datasets: I: Flight 1, II: Flight 2, III: Flight 3, IV: Flight 4, Combined: Flights I-IV merged, and for Dataset I: I0GCP: with no GCPs, I1GCP: with 1 GCP, I2GCP: with 2 GCPs, I3GCP with 3 GCPs.
Dataset Number of trees All $<$30 cm $>$30 cm RMSE (cm) All $<$30 cm $>$30 cm RMSEAll $<$30 cm $>$30 cm Bias (cm) All $<$30 cm $>$30 cm BiasAll $<$30 cm $>$30 cm Standard Deviation All $<$30 cm $>$30 cm
@@ 5.1 Autonomous flying
The performance of autonomous navigation was promising in both test forests. The obstacle avoidance success rate was 100% in the medium forest and 87.5% in the difficult forest. However, the smoothness of the trajectories suffered from the increasing forest density. While in the medium forest, 71.4% of the flights had a smooth trajectory, in the difficult forest at least one emergency stop was needed in all of the flights. Further consideration should be given to finding the optimal sensor for detecting thin and dry branches earlier. Nonetheless, the current prototype already allows reliable and safe autonomous obstacle avoidance in easy and medium difficulty level boreal forests.
The accuracy of the position estimate by VINS-Fusion was satisfying for short flights performed in this study, but in the longer flights, the error would start to accumulate making the position errors high. However, the standard deviation between the errors between the test flights was small, and the VINS-Fusion systematically estimated the flight paths shorter than in reality (as can be seen from Figure 8). This suggests, that a notable part of the position error could be eliminated by improving the calibration of the sensor setup. The results suggest that the estimates by VINS-Fusion were more accurate with online $t_{d}$ estimation than when the $t_{d}$ value was fixed to the one obtained from the calibration. However, further development of the prototype should include implementing hardware-level synchronization between the camera and the IMU of the autopilot.
The position error can also be reduced by optimizing the estimate with loop closures, and the results from the walking tests showed that VINS-Fusion could detect loop closures even in a homogeneous forest environment. Increasing the number of features tracked by VINS-Fusion increases the number of detected loops. However, increasing the number of features increases also the computational burden of the system potentially disrupting the real-time estimation. Finding the optimal value for the tracked features as well as measuring the actual improvement of the loop closures to the position estimate requires further studies.
@@ 5.2 Accuracy of forest measurements
The proposed system was assessed by comparing it to other relevant studies (Section 2.2, Table 1) considering the accuracy, system autonomy and weight, and complexity of test environment.
@@
The highest accuracy UAV SfM-MVS results were presented by Shimabuku, Konoshima, and Ota (2023) with an RMSE of 0.4 cm (3.2%). However, their test site included only 7 reference trees. Kuželka and Surovỳ (2018) and Krisanski, Taskhiri, and Turner (2020) had more reference trees and achieved similar DBH accuracy compared to that in this study, but the flights were piloted manually. Notably, the SfM-MVS approaches involved light-weight systems.
The fully automated under-canopy UAV Lidar system (ULS) by Liang et al. (2024) weighted 5 kg and provided an RMSE of 4.48 cm (21.37%) and bias of 5.01 (14.91%) at best. Wang et al. (2021) combined data from above- and under-canopy drone with a Riegl miniVUX-1UAV lidar (Riegl GmbH, Horn, Austria), resulting in an RMSE of 7.95 cm (33.35%) and bias of 3.72 cm (15.61%). Their under-canopy flight was carried out manually by a pilot and the system weighted approximately 10 kg. Muhojoki et al. (2024) carried out manually flown flights in three sparse density forest plots with the Deep Forest drone utilized an Ouster OS0-32 Rev. 5 laser scanner (Ouster, San Francisco, CA, USA) and the Hovermap drone equipped with a Velodyne VLP-16 laser scanner (Velodyne Lidar, San Jose, CA, USA), resulting in an averaged RMSE of 1.5 cm (6.7 %) and 1.6 cm (6.9 %) and bias of -0.7 cm (3.0 %) and -0.1 cm (-0.6 %). However, without utilizing their bias removal method RMSE and bias were significantly higher. The highest accuracy DBH estimates with under-canopy LiDAR were presented by Hyyppä et al. (2020a) resulting in an RMSE of 0.6 cm (2.3%) and bias of 0.34 cm (1.3%) in a sparse forest and 1.1 cm (3.5%) and 0.12 cm (0.4%) in a medium density forest. Their system was also manually piloted and weighted approximately 10 kg, a Tarot 960 UAV equipped with Kaarta Stencil-1 (VLP-16) LiDAR (Kaarta, Pittsburgh, Pennsylvania, USA).
@@
Summarizing, the highest accuracy levels were obtained with the relatively heavy ULS systems, which were operated manually in sparse forest environments. Miniaturized lightweight sensors were used in the studies by Shimabuku, Konoshima, and Ota (2023), Krisanski, Taskhiri, and Turner (2020), and Kuželka and Surovỳ (2018), and Liang et al. (2024). Results from the SfM-MVS systems were comparable to each other but ULS provided slightly poorer accuracy. Only studies by Shimabuku, Konoshima, and Ota (2023) and Liang et al. (2024) had flights with a high level of automation. Considering the complexity of the environments, the forest by Liang et al. (2024) (1000 trees/ha) was the closest to that in this study (650 and 2000 trees/ha), while other studies were conducted in remarkably sparser forests.
Based on the previous, it can be concluded that when considering the combination of accuracy, autonomy of flying, and environmental complexity, the proposed system introduced a new level of performance for under-canopy forest reconstruction. The results show the great promise of miniaturized camera SfM-MVS in autonomous under-canopy forest reconstruction.
@@ 5.3 Potential improvements and future work
During the evaluation of the system, some areas for improvement were identified, which can help develop the method towards a tool in practical forest measurements.
@@
The reliability of the navigation could be improved by performing further research on potential sensors and methods for detecting the thin branches. Blindsides in the mapping could be tackled either by sensors with 360° FOV (e.g. Lidar, 360° camera or multi-camera system) or storing historical data to occupancy grid maps in addition to the currently visible area.
@@
The virtual floor and ceiling of Ego-planner-V2 (Zhou et al. 2022) are also defined as static with respect to the initial altitude of the drone limiting the low-altitude flying to relatively flat areas. To comprehensively map uneven forest plots, the altitude limits should be modified to be dynamic and with respect to the local terrain level.
@@
The drift compensation is one of the important future development goals. Loop closure tests provided promising results and potential future tests could evaluate the approach by performing block flights. In addition to the loop closures, the error in the position estimate could potentially be further reduced by fusing GNSS data to the VIO estimate. The potential methods for GNSS-Fusion include using the global fusion features in VINS-Fusion (Qin et al. 2019a) or replacing VINS-Fusion with another method fusing GNSS-data and VIO, e.g. GVINS (Cao, Lu, and Shen 2022). Even though the dense forest canopy blocks the GNSS signals, during long-term operations the drone could receive corrections to the position estimate in the open areas while relying on the VIO estimate in denser parts of the forests.
@@
It is also notable that the rigorous photogrammetric post-processing effectively compensated for the systematic errors. While the real-time trajectory error was 50 cm, the post-processed 3D error was approximately 11 cm without any GCPs, this correlates with the trajectory error. Therefore, for applications relying on post-processing, drift is not expected to pose significant issues, especially when covering small areas such as forest test plots.
@@
In this study, the performed flight paths were straight flights through the sample plots, and the DBH estimation was performed only for the tree trunks recorded during that path. For more comprehensive forest characterization, the trajectory should cover the whole test plot, and observe the trees from multiple sides. The global path covering the whole test plot could be implemented either with a simple waypoint generation algorithm as Liang et al. (2024) or with more advanced exploration methods such as proposed by Zhou et al. (2021a) or Cieslewski, Kaufmann, and Scaramuzza (2017). Another way to achieve area coverage would be utilizing a swarm of drones as in the original article by Zhou et al. (2022). Furthermore, recent exploration algorithms (Zhou, Xu, and Shen 2023)(Bartolomei, Teixeira, and Chli 2023) are able to utilize drone swarms to further improve and speed up the area coverage.
@@
While this study only demonstrated the potential of forest characteristics estimation through DBH estimation and tree detection, the scope of data collection is not limited to that. Having a second camera pointing upwards would enable a more comprehensive reconstruction of the tree trunks to perform also stem curve estimation based on the point clouds. With a developed camera-based prototype, the data collection is not limited to point clouds, but also applications of image based mapping and object detection can be developed on top of it. Potential applications could include e.g. classifying and counting species of trees and other vegetation as well as monitoring damages in the bark of the trees due to insect pests.
@@
The prototype opens also possibilities for research of real-time forest characteristics analysis. In this study, the point clouds were obtained with post-processing, but the utilized onboard camera Intel Realsense D435 can also produce 3D point clouds in real-time. Image-based object detection can also be performed in real-time e.g. with YOLO. The analysis can be performed either with the onboard computer or with a remote computer or cloud service over the Internet connection.
6 Acknowledgements

### v2

@@ 4.2.1 Photogrammetric processing
Results of the photogrammetric processing are given in Table 5. In the image alignment process, all images were aligned in each dataset, which indicated successful processing. Re-projection errors varied between 0.29 and 0.39 pixels. The estimated GSDs (Ground Sampling Distance) ranged between 4.08 and 6.17 mm at ground level. Distances from the trees varied between 0.55 and 7.94 m, resulting in GSDs of 0.85-12.34 mm at the objects of interest. Point clouds from datasets II-IV had similar point densities of 2.63-2.84 points/cm2, while dataset I provided a slightly higher point density of 3.58 points/cm2.
Table 5: Photogrammetric processing results. Img. count: Number images/Number of aligned images; Number of GCPs; Completeness % (Number of trees found/Number of total trees); Reprojection error (pix); Number of tie points (in millions); Point density (points/cm2) in ’Evo-medium’ test site. I: Trajectory 1, II: Trajectory 2, III: Trajectory 3, IV: Trajectory 4, Combined: Trajectories 1-4 merged, and for Trajectory 1 I0GCP: no GCPs, I1GCP: 1 GCP, I2GCP: 2 GCPs, I3GCP: 3 GCPs
@@
Results of accuracy estimation using Check Points (CPs) are shown in Table 6. The cases without GCPs or with one GCP at the beginning of the trajectory provided similar 2D and 3D RMSEs of 11.36-11.72 cm and 11.37-11.73 cm, respectively. When two GCPs were used, the 2D and 3D RMSEs improved to 5.26 cm and 5.48 cm, respectively. Adding an additional GCP resulted in a significant improvement in the 2D and 3D RMSE, reducing them to 1.41 cm and 1.48 cm, respectively.
Table 6: Accuracy estimation (RMSE) using Check Points (CPs) and Camera Location in ’Evo-medium’ test site, for trajectory 1: I0GCP: no GCPs, I1GCP: 1 GCP, I2GCP: 2 GCPs, I3GCP: 3 GCPs
@@ 4.2.2 Tree detection
Figure 12 shows post-processed trajectories for dataset I-IV, as well as reference and detected trees and the boundaries of the point clouds. I and II followed almost the same trajectory, while III and IV had similar trajectories. Dataset I trajectory was the smoothest, and it did not include any sharp turns. The other three trajectories had one or more sharper turns. For datasets I, II, III, and IV, the number of reference trees inside the point cloud boundary was 29, 29, 31, and 32, of which 23, 18, 19, and 20 were found from the drone-collected point clouds, respectively. The resulting completeness values were 79.31 %, 62.07 %, 61.29 % and 62.5 %, respectively. The total number of different trees found was 29 out of 33 (87,88 %) from dataset I-IV. However, when all the flights were processed simultaneously (combined) that resulted in only $3$ completeness, where only 12 of 36 trees were found. When utilizing 0, 1, 2, and 3 GCPs in the dataset I, completeness values were 72.41 %, 68.97 %, 75.86 %, and 79.31 %, respectively (Table 5).
Figure 12: Post-processed trajectories, reference trees and found trees in Dataset (a) I (b) II (c) III (d) IV
In dataset I, the stem detection algorithm failed to detect one large tree and two small trees located near the flight trajectory despite these trees being well-reconstructed in the point cloud. Similarly, the number of well-reconstructed trees but not detected by the algorithm was 4, 5, and 5, in datasets II, III, and IV, respectively. The three extra trees that were not identified in dataset I were situated further from the trajectory and exhibited poor reconstruction quality in the point cloud. The datasets II through IV each contained seven poorly reconstructed trees. Many of these trees were located at a greater distance from the trajectory, while others were obscured from the view of the camera by intervening trees.
@@ 4.2.3 DBH estimation
RMSE and bias of the DBH estimates of datasets I-IV were similar, ranging from 3.33 to 3.86 cm (10.69 %-12.98 %) and -0.67 to 1.40 (-2.11 %-4.71 %), respectively (Table 7). Processing four flights simultaneously (Combined) provided only marginally better results with an RMSE of 2.91 cm (9.14 %) and a bias of 1.51 cm (4.74 %).
When considering the impact of the number of GCPs, the utilization of 3, 2, 1, and 0 GCPs in photogrammetric processing yielded slightly poorer results compared to those obtained with all GCPs. The RMSE and bias values ranged from 3.93 to 4.37 cm (13.68 %-14.66 %) and 1.12 to 2.13 cm (3.89 %-7.55 %), respectively (Table 7).
Tree size had an impact on the DBH estimation accuracy (Table 7). For small and medium-sized trees with a DBH smaller than 30 cm, DBH estimations achieved superior performance in all datasets, with RMSE and bias varying from 1.16 to 2.75 cm (5.74 %-12.47 %) and -0.01 to 0.62 cm (-0.06 %-3.14 %), respectively. Significantly poorer performance was obtained for trees with a DBH greater than 30 cm, with RMSE and bias ranging from 4.04 to 11.08 cm (8.99 %-25.44 %) and -0.66 to 2.08 cm (-1.53 %-5.65 %), respectively.
Analysis of the impact of the camera distance from trees did not reveal a significant correlation between tree distance and DBH estimation errors.
Table 7: Accuracy of the stem diameter breast height (DBH) in ’Evo-medium’ test site. Dataset is split in 3 categories: All, DBH $<$ 30 cm and DBH $>$ 30 cm trees. Datasets: I: Flight 1, II: Flight 2, III: Flight 3, IV: Flight 4, Combined: Flights I-IV merged, and for Dataset I: I0GCP: with no GCPs, I1GCP: with 1 GCP, I2GCP: with 2 GCPs, I3GCP with 3 GCPs.
Dataset Number of trees All $<$30 cm $>$30 cm RMSE (cm) All $<$30 cm $>$30 cm RMSE % All $<$30 cm $>$30 cm Bias (cm) All $<$30 cm $>$30 cm Bias % All $<$30 cm $>$30 cm Standard Deviation All $<$30 cm $>$30 cm
@@ 5.1 Autonomous flying
The performance of autonomous navigation was promising in both test forests. The obstacle avoidance success rate was 100 % in the ’Evo-medium’ forest and 87.5 % in the ’Evo-difficult’ forest. However, the smoothness of the trajectories suffered from the increasing forest density. While in ’Evo-medium’ 71.4 % of the flights had a smooth trajectory, in ’Evo-difficult’ at least one emergency stop was needed in all of the flights. However, apart from the one failed flight in ’Evo-difficult’, the system was able to do a successful trajectory re-planning and finish the mission after the emergency stops. Further consideration should be given to finding the optimal sensor for detecting thin and dry branches earlier. Nonetheless, the current prototype already allows reliable and safe autonomous obstacle avoidance in easy and medium-difficulty level boreal forests.
The accuracy of the position estimate by VINS-Fusion was satisfying for short flights performed in this study, but in the longer flights, the error would start to accumulate, making the position errors high. However, the standard deviation between the errors between the test flights was small, and the VINS-Fusion systematically estimated the flight paths shorter than in reality (as can be seen from Figure 9). This indicates that a notable part of the position error could be eliminated by improving the system calibration. The results indicate that the estimates by VINS-Fusion were more accurate with online $t_{d}$ estimation than when the $t_{d}$ value was fixed to the one obtained from the calibration. However, further development of the prototype should include implementing hardware-level synchronization between the camera and the IMU of the autopilot.
The position error can also be reduced by optimizing the estimate with loop closures, and the results from the walking tests show that VINS-Fusion could detect loop closures even in a homogeneous forest environment. Increasing the number of features tracked by VINS-Fusion increased the number of detected loops. However, increasing the number of features also increases the computational burden of the system, potentially disrupting the real-time estimation. Finding the optimal value for the tracked features, as well as measuring the actual improvement of the loop closures to the position estimate, requires further studies.
@@ 5.2 Accuracy of forest measurements
The proposed system was assessed by comparing it to other relevant studies (Section 2.1, Table 1) considering the accuracy, system autonomy, weight, and complexity of the test environment.
@@
The RMSEs reported in the relevant studies varied from 3.2 % to 61.41 %. The highest accuracy UAV SfM-MVS results were presented by Shimabuku, Konoshima, and Ota (2023) with RMSEs of 0.4 cm (3.2 %) and 0.7 cm (3.6 %) in two test forest sites. However, their test sites included only 7 and 18 reference trees, and the semi-autonomous commercial drone system needed manual control to avoid horizontal obstacles. Kuželka and Surovỳ (2018), He et al. (2025), and Krisanski, Taskhiri, and Turner (2020) had more reference trees and achieved similar DBH accuracy compared to that in this study, but the flights were piloted manually. Notably, the SfM-MVS approaches involved commercial lightweight drone systems, except He et al. (2025) where 360° camera was mounted as an additional payload to heavier LiDAR-based custom drone.
The most accurate under-canopy LiDAR-based DBH estimates were presented by Hyyppä et al. (2020a), achieving an RMSE of 0.6 cm (2.3 %) and bias of 0.34 cm (1.3 %) and 1.1 cm (3.5 %) and 0.12 cm (0.4 %) in two sparse forest plots. However, their system weighed around 10 kg and was piloted manually.
Autonomously operating under-canopy LiDAR-based drones were utilized in the studies by Muhojoki et al. (2024), Liang et al. (2024), and Cheng et al. (2024). While the commercial Hovermap drone system utilized in Muhojoki et al. (2024) achieved the lowest RMSE of these, the system was able to fly only in sparse and clean forests due to its size. The automated under-canopy UAV Lidar system by Liang et al. (2024) weighted 5 kg and provided an RMSE of 4.48 cm (21.37 %) and bias of 5.01 (14.91 %) at best. The autonomous system utilized in Cheng et al. (2024) weighed 4.2 kg and provided an RMSE of 3.8 cm (12.8 %) with the data collected in autonomous flight mode. The same LiDAR-based drone platform was used in studies by Prabhu et al. (2024) and He et al. (2025), but based on the articles, it seems that data collecting flights were performed in manual flight mode in these studies. Prabhu et al. (2024) reported RMSE of 1.67 cm (7.42 %) with LiDAR data, and He et al. (2025) reported median relative error of 11.5 % with photogrammetric point clouds from 360° camera data.
@@
In summary, the highest accuracy levels were obtained with the relatively heavy ULS systems operated in sparse forest environments or in environments with few reference trees. Miniaturized lightweight sensors were used in the studies by Shimabuku, Konoshima, and Ota (2023), Krisanski, Taskhiri, and Turner (2020), Kuželka and Surovỳ (2018), Prabhu et al. (2024), Cheng et al. (2024), He et al. (2025), and Liang et al. (2024), though the LiDAR-based systems (Prabhu et al. 2024; Cheng et al. 2024; He et al. 2025; Liang et al. 2024) were still considerably heavier and larger than the camera-based drone systems. From these, only studies by Shimabuku, Konoshima, and Ota (2023), Cheng et al. (2024), and Liang et al. (2024) reported flights with a high level of automation, and LiDAR-based flights in Cheng et al. (2024) and Liang et al. (2024) were the only ones where the systems were able to fly without commands from the pilot during the flight. Considering the complexity of the environments, the density of the forest in Liang et al. (2024) (1000 trees/ha) was between the forests in this study (650 and 2000 trees/ha). The natural-like subtropical forest in Shimabuku, Konoshima, and Ota (2023) had seven trees in 5 m $\times$ 5 m area, leading to a density of 2000 trees/ha, but the test plot had minimal size. In Cheng et al. (2024), the forest density was not reported.
The approach proposed in this study, camera-based autonomous under-canopy flight and photogrammetric point cloud processing from low-cost onboard stereo camera data, yielded a similar level of estimation accuracy to most of the manually piloted UAV SfM-MVS systems (Kuželka and Surovỳ 2018; Krisanski, Taskhiri, and Turner 2020; He et al. 2025). Compared to the studies with LiDAR-based autonomous UAV collection, the accuracy was higher than in Liang et al. (2024) and similar than in Cheng et al. (2024) even without any GCPs. However, the system utilized in Cheng et al. (2024) has yielded slightly better accuracy with manually piloted flights (Cheng et al. 2024; Prabhu et al. 2024), suggesting that their LiDAR-based system has the potential for higher estimation accuracy by optimizing the autonomous flight trajectories.
Although the estimation accuracy obtained in this study was in line with most of the relevant studies even without any GCPs, the reconstruction quality (completeness and RMSE) can be further improved with GCPs.
Based on the previous, it can be concluded that camera-based autonomous miniaturized drones hold the potential to become a practical tool in forest parameter estimation.The potential estimation accuracy with low-cost stereo-cameras is near to the quality of small LiDARs suitable for miniaturized drones, but the weight and the price of the sensor are considerably lower. This allows the building of smaller platforms capable of flying in denser forests, as well as reduces the cost of the system.
@@ 5.3 Potential improvements and future work
The empirical evaluation of the system revealed areas for improvement, which can inform further development toward a practical tool for operational forest measurements.
@@
The reliability of navigation and the smoothness of the trajectory could be further improved by investigating alternative camera systems and image processing methods (e.g., image super-resolution) with a better capability of detecting thin branches. Blindsides in the obstacle avoidance map could be tackled either by sensors with 360° FOV (e.g., 360° camera or multi-camera system) or storing historical data to occupancy grid maps in addition to the currently visible area.
@@
The virtual floor and ceiling of Ego-planner-V2 (Zhou et al. 2022) are also defined as static with respect to the takeoff location of the drone, limiting the low-altitude flying to relatively flat areas. To comprehensively map uneven forest plots, the altitude limits should be modified to be dynamic and concerning the local terrain level.
@@
Drift compensation remains an important objective for future development. Initial loop closure tests yielded promising results, and future developments could assess the approach using block flights with a lawnmower pattern. In addition to the loop closures, the error in the position estimate could potentially be further reduced by fusing GNSS data to the VIO estimate. The potential methods for GNSS-Fusion include using the global fusion features in VINS-Fusion (Qin et al. 2019a) or replacing VINS-Fusion with another method fusing GNSS-data and VIO, e.g. GVINS (Cao, Lu, and Shen 2022). Even though the dense forest canopy blocks the GNSS signals, during long-term operations, the drone could receive corrections to the position estimate in the open areas while relying on the VIO estimate in denser parts of the forests.
@@
It is also notable that the rigorous photogrammetric post-processing effectively compensated for the systematic errors. While the real-time trajectory error was 50 cm, the post-processed 3D error was approximately 11 cm without any GCPs.Therefore, for applications relying on post-processing, drift is unlikely to present significant challenges.
@@
In this study, the performed flight paths were straight flights through the sample plots, and the DBH estimation was performed only for the tree trunks recorded during that path. For more comprehensive forest characterization, the trajectory should cover the whole test plot and observe the trees from multiple sides. Reconstructing the tree trunk from multiple sides instead of only one side could possibly also reduce the higher DBH estimation error obtained with the trees with large diameters. The global path covering the whole test plot could be implemented either with a simple waypoint generation algorithm as Liang et al. (2024) or with more advanced exploration methods such as proposed by Zhou et al. (2021a) or Cieslewski, Kaufmann, and Scaramuzza (2017). Another way to achieve area coverage would be utilizing a swarm of drones as in the original article by Zhou et al. (2022). Furthermore, recent exploration algorithms (Zhou, Xu, and Shen 2023)(Bartolomei, Teixeira, and Chli 2023) are able to utilize drone swarms to further improve and speed up the area coverage.
@@
While this study focused on demonstrating the potential for forest characteristics estimation through DBH estimation and tree detection, the scope of data collection extends well beyond these applications. Including an additional lightweight camera pointing upwards would enable a more comprehensive reconstruction of the tree trunks to perform also stem curve estimation based on the point clouds. With the developed camera-based prototype, data collection extends beyond point clouds to encompass applications such as image-based mapping and object detection. This opens up a variety of potential uses, including classifying and counting tree species and other vegetation, as well as monitoring bark damages caused by insect pests.
@@
The prototype also opens possibilities for research of real-time forest characteristics analysis. In this study, the point clouds were obtained through post-processing. However, the onboard Intel Realsense D435 camera is also capable of generating 3D point clouds in real time. Additionally, image-based object detection can be performed in real-time, e.g., with YOLO (Redmon et al. 2016). While only computationally light analysis can be conducted on the onboard computer, real-time analysis with higher computational burden could be achieved through a remote computer or cloud service via an Internet connection.
Acknowledgements